Fermi National Laboratory
 

Accelerator Update for October 6, 2006 - October 9, 2006

 

Friday October 6

The day shift began with Operations monitoring the Recycler stash (144.8E10) and the Antiproton Source (Pbar) stack (50.9E10), with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the colliding beam, with the Tevatron (TeV) in shot setup, with Linac, Booster, Main Injector (MI) operating normally, and with NuMI and MiniBooNE taking beam.

At 10:13 AM, Operations reported that a P1 transfer line power supply went too high and caused the line’s beam permit to trip followed by the NuMI permit. This occurred several times during the day shift.

Operations established store 5006 at 10:55 AM with an initial average luminosity of 114.2E30.

A Recycler expert began an hour-long electron cooling study at 10:56 AM.

Operations reported Pbar controls issues at 12:02 PM. Experts investigated.

The Recycler completed stashing antiprotons at 2:11 PM.

A MI expert began a multi-batch slipstacking study at 2:14 PM.

Operations reported at 6:40 PM a New Pbar Stacking Record with a rate of 20.56mA per hour.

The Recycler completed a stash at 7:41 PM.

At 8:31 PM, the NuMI horn tripped off. It reset okay.

At 8:39 PM, a Booster RF station (BRF6) tripped off on a power amplifier LCW flow fault. It would not reset. Operators bypassed the station.

The Recycler completed a stash at 11:42 PM.

Saturday October 7

The midnight shift began with Operations monitoring store 5006 (luminosity 35.37E30) and the stash (234.2E10), with Pbar stacking (216.8E10), and with NuMI and MiniBooNE taking beam.

The Recycler completed a stash at 1:41 AM.

At 2:13 AM, an Accumulator RF station (ARF1) tripped off. Operators noted a burning smell and called experts.

At 3:25 AM, the Duty Mechanic reported a possible LCW leak in Pbar. Operators prepared the area for an access.

At 3:44 AM, operators found the LCW leak on the ARF1-2 cavity. Operators valved off the LCW and called in experts.

The Recycler completed a stash at 7:19 AM.

At 8:26 AM, the H- Source extraction voltage dropped, which caused the Linac beam intensity to drop. The voltage slowly came back to its normal level.

The Recycler completed a stash at 9:17 AM.

The Recycler completed a stash at 11:25 AM.

Operations terminated store 5006 at 11:50 AM.

Operations put the TeV into shot setup at 12:07 PM.

Operators prepared Booster for an access at 2:07 PM to check BRF6.

Operations established store 5007 at 2:56 PM with an initial luminosity of 192.58E30.

Booster experts reported at 3:17 PM that they had replaced the BRF6 flow meter, but that didn’t fix the problem. During circuit diagnostics, the problem went away. Experts could not make the system fail. Operations turned Booster back on at 3:53 PM.

The Recycler completed a stash at 5:37 PM.

At 5:52 PM, operators mistakenly changed the TeV proton tune instead of the antiproton tune. This caused a loss of luminosity.

The Recycler completed a stash at 7:42 PM.

Operations reported at 8 PM that store 5007’s luminosity was at 81.87E30, the stack was at 20.3E10, and the stash was at 83.7E10.

At 10:18 PM, a Booster kicker (MKS07) tripped off. Operators believed its cable needed to be reterminated and called in an expert.

Sunday October 8

The midnight shift began with Operations monitoring store 5007 and the stash, but with Pbar, NuMI, and MiniBooNE not taking beam due to a Booster kicker problem.

A Booster expert repaired MKS07 and Booster resumed sending beam at 12:19 AM.

The Recycler completed a stash around 1:24 AM.

The Recycler completed a stash at 5:05 AM.

The Recycler completed a stash at 8:15 AM.

The Recycler completed a stash at 11:16 AM.

Experts began a NuMI/Pbar interleaving study at 1:19 PM.

The Recycler completed a stash at 3:22 PM.

At 3:26 PM, Operations reported that the H- Source had suffered some sparking, which reduced intensity. It recovered slowly.

The Recycler completed a stash at 6:42 PM.

TeV experts began an end-of-store study at 6:36 PM.

Operations reported at 8 PM that the stack was 37.3E10 and the stash was at 376.9E10.

Operations terminated store 5007 at 8:41 PM and then put the TeV into shot setup about ten minutes later.

Operations established store 5008 at 11:12 PM with an initial luminosity of 206.95E30.

Operations prepared the Recycler to stash at 11:27 PM, but problems with ARF4 messed up three transfer attempts. No antiprotons were lost due to this problem. Operators called in and expert.

Monday October 9

The midnight shift began with Operations monitoring store 5008 and the stash, with Pbar stacking, with a RF expert investigating an ARF4 problem, and with NuMI and MiniBooNE taking beam.

The Recycler completed this problematic stash at 12:12 AM.

With an expert monitoring ARF4, the Recycler successfully completed another stash at 1:53 AM.

At 2:51 AM, Operations reported a problem with store 5008’s luminosity lifetime. Operators adjusted the tunes.

The Recycler completed a stash at 5:49 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinators will allow TeV experts to conduct a Tevatron Electron Lens end-of-store study.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/09/2006    email Fermilab